# Unlocking Doping Effects on Altermagnetism in MnTe

### v1

@@ III.2 Magnetic structure of MnTe with single non-magnetic atom substitution
III.4 Anomalous Hall effect in MnTe with paired non-magnetic atom substitutions
@@
The existence of AHC in altermagnets is not inherently guaranteed and depends on the orientation of the Néel vector. In hexagonal MnTe, when the magnetization is aligned along the $c$-axis, AHC is suppressed due to the presence of symmetry operations such as $C^{\pm}_{3z}$ and in-plane two-fold rotations $C_{210}$ and $C_{120}$ within the MSG $P6_{3}^{\prime}/m^{\prime}m^{\prime}c$ (No.194.268) [21]. In this section, we investigate the potential for generating a non-zero AHC in MnTe through doping-induced symmetry breaking, based on general symmetry analysis. For a single dopant substitution, or a pair atom substitution with MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211), the symmetries that prohibit AHC are retained. However, for other pair doping configurations, with reduced MSG $P\bar{3}m^{\prime}1$ (No.164.89), a Hall pseudovector along the $c$-axis becomes symmetry-allowed, as the orthogonal rotation axes combined with time-reversal symmetry ($\tau C_{2x}$, $\tau C_{2y}$) leave it invariant. In lower-symmetry structures with MSGs $Fm^{\prime}m^{\prime}2$ (No.42.222) and $Pm^{\prime}m^{\prime}2$ (No.25.60), the presence of the $C_{2z}$ symmetry operation enforces the vanishing of AHC components $\sigma_{yz}$ and $\sigma_{xz}$, while imposing no constraints on $\sigma_{xy}$. On the other hand, the lowest-symmetry and the most probable configuration with MSG $C2^{\prime}/m^{\prime}2$ (No.12.62) lacks any pure rotational symmetry and retains only the combined symmetry operation $\tau C_{2y}$, which permits non-zero components $\sigma_{yz}$ and $\sigma_{xy}$.
While changing the dopant position offers the possibility to tune the direction of the AHC, changing the type of dopant provides a way to tune the magnitude of the AHC near the Fermi level. For instance, Se atoms, which have the same number of valence electrons as Te, preserve the large band gap of MnTe, resulting in a vanishing AHC near the Fermi level. In contrast, doping with Sb or I shifts the Fermi level toward the valence or conduction bands, respectively, leading to finite AHC values at the Fermi level. In Fig. 5 we show the non-zero components of AHC for candidate materials in two distinct classes of pair non-magnetic atom doped MnTe systems – an altermagnet and a quasi-altermagnet – each incorporating Sb and I atoms as dopants. Top panel illustrates $\sigma_{xy}$ component in the altermagnetic case with $Pm^{\prime}m^{\prime}2$ (No.25.60), where the left (right) one corresponds to Sb (I) doped systems. In the bottom panel, we present two non-zero AHC components – $\sigma_{xy}$ and $\sigma_{yz}$ – for quasi-altermagnets with MSG $C2^{\prime}/m^{\prime}2$ (No.12.62). Our findings are in agreement with the symmetry considerations discussed previously.
@@
III.5 Microscopic origin of altermagnetism in doped MnTe: A model Hamiltonian approach
@@
To gain insights into the microscopic origin of formation and dissipation of altermagnetism in doped MnTe, we develop a minimal tight-binding (TB) Hamiltonian based on Slater-Koster (SK) formalism. The SKTB Hamiltonian that exhibits altermagnetism in this class of materials is given by [25],
@@
where, i,ji,\ j are site indices, $\mu,\ \nu$ depict orbital indices, and $\sigma$ describes the spin indices. $\tau=1$ represents the sublattice where spin-up states form the majority spin channel, and $\tau=2$ represents the sublattice where spin-down states form the majority spin channel. The detailed analysis for the chemical bonding origin of altermagnetism in NiAs prototype compounds can be found in Ref. [25]. We primarily focus on the pair doping cases, as single doping is a subset of this. Here, we intend to see the effect of doping in MnTe, which will affect only a part of the existing chemical bonding of the pristine crystal. Since the objective here is to get the underlying mechanism responsible for the presence or absence of altermagnetism and not the full band structure, we created a few quasi-1D crystals as models to study the role of doped atom driven chemical bonding. For further simplicity, we adapted a single orbital model (i.e., one $d$-orbital from the magnetic atom and one $p$-orbital from the non-magnetic atom), as the intra-orbital interactions do not have a significant role in the formation of altermagnetism. Here, individual magnetic atoms form the basis for the magnetic sites, while a linear combination of three nearest-neighbor non-magnetic atoms of a given magnetic atom in the same $xy$-plane form the basis for the non-magnetic sites. We take $d_{xy}$ orbital of the magnetic atom and $p_{x}$ orbital for the non-magnetic atom as an example. A similar analysis can be carried out for all other combinations of the orbitals in an analogous manner. Let us next consider the two cases with altermagnetism and quasi-altermagnetism.
Figure 6: SKTB analysis of altermagnetism in MnTe with pair doping. The primitive unit cell of the representative crystal of MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211) [see Fig. 3(a)] is considered here. Panels (a)-(c) represent the three quasi-1D configurations (opaque color), which include the doped atoms. The nearest-neighbor non-magnetic coordination of a given magnetic atom is shown by the triangular plane. In (b) and (c), some additional nonmagnetic atoms are generated by employing periodic boundary conditions to represent the quasi-1D chains properly. Panel (d) shows the band structure of configuration 1, where the altermagnetism is not affected by doping, with the expected spin-splitting retained.
@@
Case I: Altermagnetism– Out of five different magnetic space group families of pair doped configurations, $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211), $Fm^{\prime}m^{\prime}2$ (No.42.222), and $Pm^{\prime}m^{\prime}2$ (No.25.60) come under this category, where a cell exists with half the volume containing only one doped atom. One representative configuration from each of these families is shown in Fig. 3(a). We consider the crystal structure representing the MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211) [see Fig. 3(a)] as a prototype to study the effects of chemical doping in this category. Here the doped atoms are placed on top of each other along $z$ and in alternate $xy$-planes. Therefore, it is possible to build a smaller supercell by reducing the size from $2\times 2\times 2$ to $2\times 2\times 1$ in this case. As discussed earlier, in the new supercell, there exist four opposite spin pairs of magnetic atoms, and each pair creates a quasi-1D configuration. Out of four such configurations, the nearest neighbor coordination of one configuration is not affected by the doping, and the corresponding chemical bonding nature remains unchanged. The other three configurations are shown in Fig. 6. The Hamiltonian for these three configurations is written as
@@
The basis order in the above Hamiltonian is $|\text{Mn}-d\rangle,\ |\text{Mn}^{\prime}-d\rangle,\ |\text{NM}_{\text{eff}}-p\rangle,\ |\text{NM}^{\prime}_{\text{eff}}-p\rangle$. Here, $|\text{NM}_{\text{eff}}-p\rangle$ is formed out of all Te atoms, while $|\text{NM}^{\prime}_{\text{eff}}-p\rangle$ includes one doped atom [see Fig. 6(a-c)]. From the figure, we see that the coordination of the up-spin magnetic atoms is exactly the mirror image of that of the down-spin magnetic atoms. Therefore, the interactions among the magnetic and non-magnetic atoms satisfy the following relations: $w_{1}^{2}+w_{2}^{2}=v_{1}^{2}+v_{2}^{2}$ and $w_{3}^{2}+w_{4}^{2}=v_{3}^{2}+v_{4}^{2}$. As a result, altermagnetism survives in this case even after chemical doping, as clearly seen in Fig. 6(d).
Figure 7: SKTB analysis of quasi-altermagnetism in MnTe with pair doping. The quasi-1D chains are designed out of the representative crystal of MSG $C2^{\prime}/m^{\prime}$ (No.12.62) as depicted in Fig. 3(b) to study the effect of doping. Panels (a) and (b) represent the two quasi 1D chains (opaque color) which include the non-magnetic doped atoms. Here also the nearest-neighbor nonmagnetic coordination of a given magnetic atom is shown by the triangular plane, and in (b), a few additional nonmagnetic atoms are generated via the lattice translation vectors to represent the quasi-1D chains properly. Panel (c) showss the altermagnetic band structure for configuration 1 without doping. Panels (d), (e), and (f) represent the quasi-altermagnetic band structure for configuration 1 due to the difference in non-magnetic onsite energies, hopping strengths, and magnetic moments, respectively.
Case II: Quasi-altermagnetism– As we discussed previously, chemical doping at any other site except the above-mentioned cases, resulted in two magnetic space group families, namely $C2^{\prime}/m^{\prime}$ (No.12.62) and $P\bar{3}m^{\prime}1$ (No.164.89). In these cases, the ideal band dispersion defining altermagnetism disappears [see Fig. 3(b) and (d)]. To investigate the deviation from ideal altermagnetic band dispersion, we take the representative crystal structure of MSG $C2^{\prime}/m^{\prime}2$ (No.12.62) as a prototype. This system can be divided into four quasi-1D configurations. The coordination of one-fourth of the magnetic atoms in the unit cell is free from doped atoms forming one quasi-1D chain, and thereby, the corresponding chemical bonding remains unaffected. The band structure in these cases can be obtained from the single-band model of pristine MnTe [25]. For another one-fourth of magnetic atoms, which form another quasi-1D chain, the coordination of opposite-spin magnetic atoms can be connected pairwise via a mirror, and the corresponding band structure can be obtained from Eq. 2. These two chains interpenetrate each other and occupy half of the magnetic sites, and the bands formed by them are altermagnetic in nature. The remaining magnetic atoms form the other two quasi-1D chains [see Fig. 7(a) and (b)].
The Hamiltonian describing their band structure can be written as
@@
Here, the basis order is $|\text{Mn}_{1}-d\rangle,\ |\text{Mn}_{2}-d\rangle,\ |\text{Mn}_{3}-d\rangle,\ |\text{Mn}_{4}-d\rangle,\ |\text{NM}_{\text{eff}1}-p\rangle,\ |\text{NM}_{\text{eff}2}-p\rangle,\ |\text{NM}_{\text{eff}3}-p\rangle,\ |\text{NM}_{\text{eff}4}-p\rangle$. Each of the non-magnetic effective bases $|\text{NM}_{\text{eff}1}-p\rangle,\ |\text{NM}_{\text{eff}2}-p\rangle$ contains one doped atom, whereas there are no doped atoms in the other two non-magnetic effective bases. The undoped chain exhibits a perfect altermagnetic band structure, as seen in Fig. 7(c). The doped atoms introduce additional perturbations to the non-magnetic onsite energy ($\beta_{1}\neq\beta_{2}$), hopping strengths ($w_{1}+iw_{2}\neq w_{7}+iw_{8}$, $v_{1}+iv_{2}\neq w_{5}+iw_{6}$, etc.), and magnetic moments ($\Delta_{\text{Mn}}\neq\Delta^{\prime}_{\text{Mn}}$). Each perturbation independently lifts the antiferromagnetic sublattice band degeneracy. In the case of isovalent doping, such as with Se, the difference between $\beta_{1}$ and $\beta_{2}$ is much more prominent than the other two perturbations. In the case of hole-type (Sb) or electron-type (I) doping, each of the perturbations become prominent leading to larger deviation from the ideal altermagnetic band structure, which is substantiated through Fig. 7(d-f). From Fig. 7(d), we see that the difference in the onsite energy of non-magnetic atoms primarily lifts the spin degeneracy of the bands at the $\Gamma$ point, which is also observed in our DFT-obtained bands, as shown in Fig. 3(b). For the non-isovalent doping cases (i.e., Sb/I doping), the uncompensated magnetic moments further strengthen the breaking of the degeneracy with larger splitting at the $\Gamma$ point [see Fig. 4(b) and (d) and Fig. 7(f)]. In addition to this, the non-isovalent doping substantially changes the strengths of hopping interactions, which introduce additional anisotropy to the spin density. Therefore, the bands deviate significantly from the ideal altermagnetic band structure as shown in Fig. 7(e). This is also reflected in our DFT band structures [see Fig. 4(b) and (d)]. Overall, our SKTB models yield a microscopic understanding of the origin of altermagnetism and quasi-altermagnetism in doped MnTe.

### v2

@@ III.4 Quasi-altermagnetism
III.5 Anomalous Hall effect in MnTe with paired non-magnetic atom substitutions
The anomalous Hall effect (AHE) is a phenomenon that arises due to broken TRS and strong SOC. Altermagnets having strong SOC can exhibit AHC as the TRS is broken. However, the existence of AHE in altermagnets is not inherently guaranteed since it crucially depends on the orientation of the Néel vector. Hexagonal MnTe is one of the prominent altermagnetic materials in which AHE is both theoretically predicted and experimentally observed [19, 21]. But this is only when the Néel vector lies in the basal plane. When Néel vector is oriented along the out-of-plane direction, pristine MnTe does not exhibit AHE. In this configuration, AHC is symmetry-forbidden due to the presence of operations such as the three-fold rotations $C^{\pm}_{3z}$ and in-plane two-fold rotations $C_{210}$ and $C_{120}$ consistent with MSG $P6_{3}^{\prime}/m^{\prime}m^{\prime}c$ (No. 194.268) [21].
@@ III.5 Anomalous Hall effect in MnTe with paired non-magnetic atom substitutions
In this section, we investigate the potential for generating a non-zero AHC in MnTe with Néel vector along the out-of-plane direction, through doping induced symmetry breaking. Similar to the pristine case, MnTe with either a single (non-magnetic) substitution or a pair (non-magnetic) substitution configuration with MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211) does not exhibit AHE. It is because the symmetry operations that prohibit AHE remain preserved in these configurations. However, in the other pair doping configurations with reduced symmetry, AHE becomes symmetry allowed, leading to a finite AHE. For the pair doped configuration with MSG $P\bar{3}m^{\prime}1$ (No.164.89), a Hall pseudovector along the $c$-axis becomes symmetry-allowed, as the orthogonal rotation axes combined with TRS ($\tau C_{2x}$, $\tau C_{2y}$) leave it invariant. In lower-symmetry structures with MSGs $Fm^{\prime}m^{\prime}2$ (No.42.222) and $Pm^{\prime}m^{\prime}2$ (No.25.60), the presence of the $C_{2z}$ symmetry operation enforces the vanishing of AHC components $\sigma_{yz}$ and $\sigma_{xz}$, while imposing no constraints on $\sigma_{xy}$. On the other hand, the lowest-symmetry and the most probable configuration with MSG $C2^{\prime}/m^{\prime}2$ (No.12.62) lacks any pure rotational symmetry and retains only the combined symmetry operation $\tau C_{2y}$, which permits non-zero components $\sigma_{yz}$ and $\sigma_{xy}$.
The intrinsic AHE originates from a non-zero Berry curvature ($\Omega$) in momentum space [45]. Within the framework of linear response theory, $\gamma\delta$-th ($\gamma\neq\delta$) component of the Hall conductivity tensor, $\sigma$, can be written in terms of the $\nu$-th component of $\Omega$ as, [45]
$\sigma_{\gamma\delta}=\varepsilon_{\gamma\delta\nu}\frac{e^{2}}{\hbar}\sum_{n}\int_{BZ}f(\epsilon_{n}(\textbf{k}))\Omega_{n}^{\nu}(\textbf{k})\frac{d\textbf{k}}{(2\pi)^{3}}.$ (1)
Here $e$, $\hbar$, $n$, $\varepsilon$, and $f(\epsilon_{n}(\textbf{k}))$ represent the electron charge, reduced Planck’s constant, Bloch band index, Levi-Civita symbol, and the Fermi-Dirac distribution function, respectively. Therefore, when the Berry curvature component is compensated over the entire Brillouin zone, the corresponding AHC component vanishes.
In Fig. 7, we show the non-zero components of AHC and $\Omega$ for candidate materials in two distinct classes of pair non-magnetic atom doped MnTe systems – an altermagnet and a quasi-altermagnet – each incorporating Sb and I atoms as dopants. From the figure, we find that for MSG 25.60 (both in Sb-doped [in (e)] and I-doped [in (g)] cases), the net contribution coming from $\Omega_{x}$ and $\Omega_{y}$ vanishes, as there exists always a negative counterpart for each of the positive Berry curvature. In contrast, $\Omega_{z}$ component has a net contribution over the entire Brillouin zone–contributing to non-zero $\sigma_{xy}$ as obtained from equation 1.
In a similar way, for MSG. 12.62, in Fig. 7 (f) and (i), one can note that, only $\Omega_{y}$ component gets compensated, and other two ($\Omega_{x}$ and $\Omega_{z}$) remain uncompensated throughout the Brillouin zone. This gives rise to two non zero components of AHC – $\sigma_{yz}$ and $\sigma_{xy}$ – as depicted in Fig. 7 (b) and (d), respectively.
These results demonstrate that non-magnetic atom doping can induce AHE in MnTe while retaining the magnetization along the out-of-plane direction, where AHC is symmetry-forbidden in the pristine case. By altering the relative dopant position, and thereby modifying the symmetry of the MnTe, it becomes possible not only to induce AHC but also to vary its direction. Furthermore changing the type of dopant provides a way to tune the magnitude of the AHC near the Fermi level. For instance, Se atoms, which are isovalent with Te, preserve the large band gap of MnTe, resulting in a vanishing AHC near the Fermi level. In contrast, doping with hole rich Sb or electron rich I shifts the Fermi level toward the valence or conduction bands, respectively, leading to finite AHC values at the Fermi level. Our work distinctly demonstrates the doping induced emergence of AHE in MnTe, both in the altermagnetic and quasi-altermagnetic configurations, while Néel vector orientation is retained in the out-of-plane direction.
@@
III.6 Microscopic origin of altermagnetism in doped MnTe: A model Hamiltonian approach
@@
To gain insights into the microscopic origin of formation and dissipation of altermagnetism in doped MnTe, we develop a minimal tight-binding (TB) Hamiltonian based on Slater-Koster (SK) formalism. The SK-TB Hamiltonian that exhibits altermagnetism in this class of materials is given by [25],
@@
where, i,ji,\ j are site indices, $\mu,\ \nu$ depict orbital indices, and $\sigma$ describes the spin indices. $\tau=1$ represents the sublattice where spin-up states form the majority spin channel, and $\tau=2$ represents the sublattice where spin-down states form the majority spin channel. The detailed analysis for the chemical bonding origin of altermagnetism in NiAs prototype compounds can be found in Ref. [25]. We primarily focus on the pair doping cases, as single doping is a subset of this. Here, we intend to see the effect of doping in MnTe, which will affect only a part of the existing chemical bonding of the pristine crystal.
If we examine the crystal structure of MnTe (for all NiAs prototypes) it can be presented as a series of chainlike interaction blocks, as shown in Figs. 8(a) and 9(a). In the figures, each coloured cuboid within the unit cell represents one such interaction block. If the electron-electron hopping interactions are restricted to the second nearest neighbour, these blocks are isolated from each other to behave like a ”quasi-1D” chain. Therefore, the up- and down-spin blocks of the total Hamiltonian matrix can be written in terms of a set of block diagonal matrices with each block representing one of this quasi-1D chain. We note that, by ignoring the interaction beyond the second neighbour the altermagnetic characteristics are not affected  [25]. As the sole objective is to demonstrate how altermagnetic band dispersion gets affected in the presence of doping, for simplicity, we adapted a single orbital model (i.e., one $d$-orbital from the magnetic atom and one $p$-orbital from the non-magnetic atom). Earlier we have exclusively shown that it is possible to demonstrate altermagnetism using a single orbital picture. Within this framework, the basis of the Hamiltonian is now formed by the two $d$-orbitals, one each from the opposite spin sublattices, representing the magnetic site, and two effective hybridized orbitals representing the non-magnetic site. There are three nearest neighbour non-magnetic isoplanar atoms forming a triangle, one below a magnetic site and one above it [see right panel Fig. 8 (a)]. For each of the triangles, one effective hybridized orbital is formed out of the linear combination of three identical $p$-orbitals, coming from the non-magnetic atoms.
We take the $d_{xy}$ orbital of the magnetic atom and the $p_{x}$ orbital for the non-magnetic atom as a case study. A similar analysis can be carried out for all other combinations of the orbitals in an analogous manner. Let us next consider the two cases with altermagnetism and quasi-altermagnetism.
Figure 8: SK-TB analysis of altermagnetism in MnTe with pair doping. The primitive unit cell of the representative crystal of MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211) [see Fig. 3(a)] is considered here. (a) Identification of different quasi-1D interaction blocks shown by colour cuboids. From the electronic structure point of view, they are isolated from each other, if the hopping interactions are restricted up to the second nearest neighbours. In each quasi-1D block, a magnetic atom is capped by two iso-planar triangles, one above and one below, formed by the non-magnetic atoms. For the single-doped case, there exist four such quasi-1D interaction blocks, out of which one remains unaffected from doping (shown in grey) and the other three contain one doped atom Se/Sb/I. (b) Schematic illustration of the block diagonal tight-binding model Hamiltonian with each block representing one of the quasi-1D chains. (c) A visualization of the hopping interactions in the Hamiltonian matrix of one of the quasi-1D chains. (d) The band structure of configuration (c) is shown, where the altermagnetism is not affected by doping, with the characteristic AMSS retained.
Figure 9: SK-TB analysis of quasi-altermagnetism in MnTe with pair doping. The quasi-1D interaction blocks are designed for the representative crystal of MSG $C2^{\prime}/m^{\prime}$ (No.12.62) as depicted in Fig. 3(b) to study the effect of doping. (a) There exist four different interaction blocks in the primitive unit cell in the framework of up to second neighbour interactions, which are shown by the colour cuboids. Out of them, two interaction blocks contain only one doped atom (as shown by green and magenta colours in the upper panel), and the remaining two interaction blocks contain two doped atoms (as shown by brown and blue colours in the middle panel). While in the former case, the altermagnetic band structure is retained due to the presence of mirror symmetry, it vanishes, and a quasi-altermagnetic band structure is formed in the latter case because of the breaking of mirror symmetry. (b) Schematic illustration of the block diagonal tight-binding model Hamiltonian, with each block representing one of the quasi-1D chains. (c) The various hopping interactions are presented for a pair doped interaction block. Quasi-altermagnetic band structure for this configuration when there is a difference between the nonmagnetic onsite energies (d), hopping strengths (e), and local spin moments (f). In a real system, all three of them contribute together.
@@
Case I: Altermagnetism– Out of five different MSG families of pair doped configurations, $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211), $Fm^{\prime}m^{\prime}2$ (No.42.222), and $Pm^{\prime}m^{\prime}2$ (No.25.60) come under the altermagnetism category, where a primitive unit cell exists with half the volume of the supercell and it contains only one doped atom. We consider the crystal structure representing MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211) [see Fig. 3(a)] as a prototype to study the effects of chemical doping in this category. Here, the doped atoms are placed on top of each other along $z$ and in alternate $xy$-planes. Therefore, in this case, it is possible to build the primitive unit cell by reducing the size of the supercell from $2\times 2\times 2$ to $2\times 2\times 1$. In this primitive unit cell, there exist four opposite spin pairs of magnetic atoms, and each pair creates a quasi-1D interaction block as shown by coloured cuboids in Fig.  8(a). Out of four such blocks, the nearest neighbor coordination of one block is not affected by the doping, and the corresponding chemical bonding nature remains unchanged (see the grey coloured cuboid), while the chemical bondings of other three blocks are affected by the doped atom (see the magenta, orange, and green blocks). The Hamiltonian for these three configurations is written as
@@
The basis order in the above Hamiltonian is $|\text{Mn}-d\rangle,\ |\text{Mn}^{\prime}-d\rangle,\ |\text{NM}_{\text{eff}}-p\rangle,\ |\text{NM}^{\prime}_{\text{eff}}-p\rangle$. Here, $|\text{NM}_{\text{eff}}-p\rangle$ is formed out of all Te atoms, while $|\text{NM}^{\prime}_{\text{eff}}-p\rangle$ includes one doped atom [see Fig. 8(a)]. The various hopping interaction terms in the Hamiltonian for the orange coloured interaction block are shown in Fig. 8(b) as an example. From the figure, we see that the coordination of the up-spin magnetic atoms is exactly the mirror image of that of the down-spin magnetic atoms. Therefore, the interactions among the magnetic and non-magnetic atoms obey the following relations: $w_{1}^{2}+w_{2}^{2}=v_{1}^{2}+v_{2}^{2}$ and $w_{3}^{2}+w_{4}^{2}=v_{3}^{2}+v_{4}^{2}$, and these relations lead to altermagnetic band structure [25], which is further validated by the band dispersions shown in Fig. 8(c).
Case II: Quasi-altermagnetism– As we discussed previously, chemical doping at any other site except the above-mentioned cases, resulted in two MSG families, namely $C2^{\prime}/m^{\prime}$ (No.12.62) and $P\bar{3}m^{\prime}1$ (No.164.89). In these cases, the ideal band dispersion defining altermagnetism disappears [see Fig. 3(b) and (d)]. To investigate the deviation from ideal altermagnetic band dispersion, we take the representative crystal structure of MSG $C2^{\prime}/m^{\prime}2$ (No.12.62) as a prototype. This system can be divided into four quasi-1D interaction blocks. Out of them, two are single doped and other two are pair doped. For the single doped interaction blocks, the characteristic AMSS is retained as discussed in the previous case.
For the pair doped quasi-1D chains, the Hamiltonian takes the shape:
@@
Here, the basis order is $|\text{Mn}_{1}-d\rangle,\ |\text{Mn}_{2}-d\rangle,\ |\text{Mn}_{3}-d\rangle,\ |\text{Mn}_{4}-d\rangle,\ |\text{NM}_{\text{eff}1}-p\rangle,\ |\text{NM}_{\text{eff}2}-p\rangle,\ |\text{NM}_{\text{eff}3}-p\rangle,\ |\text{NM}_{\text{eff}4}-p\rangle$. The triangular planes forming the effective orbitals $|\text{NM}_{\text{eff}1}-p\rangle,\ |\text{NM}_{\text{eff}2}-p\rangle$ have one doped atom. The other tringular planes forming the effective orbitals $|\text{NM}_{\text{eff}3}-p\rangle,\ |\text{NM}_{\text{eff}4}-p\rangle$ do not have any doped atoms. In Fig. 9(b), various hopping interactions in the Hamiltonian of Eq. 4 are schematically shown for the brown block as an example. The pair doped atoms introduce perturbations to the non-magnetic onsite energy ($\beta_{1}\neq\beta_{2}$), hopping strengths ($w_{1}+iw_{2}\neq w_{7}+iw_{8}$, $v_{1}+iv_{2}\neq w_{5}+iw_{6}$, etc.), and magnetic moments ($\Delta_{\text{Mn}}\neq\Delta^{\prime}_{\text{Mn}}$). Each perturbation independently lifts the antiferromagnetic sublattice band degeneracy. In the case of isovalent doping, such as with Se, the difference between $\beta_{1}$ and $\beta_{2}$ is much more prominent than the other two perturbations. In the case of hole-type (Sb) or electron-type (I) doping, each of the perturbations becomes prominent, leading to a larger deviation from the ideal altermagnetic band structure, which is substantiated through Fig. 9(d-f).
From Fig. 9(d), we see that the difference in the onsite energy of non-magnetic atoms primarily lifts the spin degeneracy of the bands at the $\Gamma$ point, which is also observed in our DFT-obtained bands, as shown in Fig. 3(b). For the non-isovalent doping cases (i.e., Sb/I doping), the uncompensated magnetic moments further strengthen the breaking of the degeneracy with larger splitting at the $\Gamma$ point [see Fig. 4(b) and (d) and Fig. 9(f)]. In addition to this, the non-isovalent doping substantially changes the strengths of hopping interactions, which introduce additional anisotropy to the spin density. Therefore, the bands deviate significantly from the ideal altermagnetic band structure as shown in Fig. 9(e). This is also reflected in our DFT band structures [see Fig. 4(b) and (d)]. Overall, our SKTB models yield a microscopic understanding of the origin of altermagnetism and quasi-altermagnetism in doped MnTe.